OpenStudy (mallorysipp234):

i'm trying to understand my final practice on paper for my final How do i go about doing this problem? Its in the comments thank you

OpenStudy (mallorysipp234):

\[\log _{3}(x)+\log_{3}(x-4)=2\]

OpenStudy (mallorysipp234):

sorry that's supposed to be subtraction not addition

OpenStudy (agent0smith):

log(a) + log(b) = log(a*b)

OpenStudy (mallorysipp234):

then it would be division

OpenStudy (agent0smith):

log(a) - log(b) = log(a/b)

OpenStudy (mallorysipp234):

\[\log _{3}(x \div x-4)=2\]

OpenStudy (agent0smith):

\[\Large \log_a b = c\]then \[\Large b = a^c\]

OpenStudy (mallorysipp234):

now i just simply solve?

OpenStudy (mallorysipp234):

my answer sheet is telling me i need to divide by 8, i don't know where the 8 came from

OpenStudy (mallorysipp234):

@agent0smith

OpenStudy (agent0smith):

Did you see my last post...?

OpenStudy (mallorysipp234):

@agent0smith

OpenStudy (mallorysipp234):